OpenStudy (anonymous):

a _{n}=((-1)^{n-1}(6)+2)/(2n+1+(-1)^{n-1}) find limit as n goes to infinity

OpenStudy (anonymous):

\[a _{n}=((-1)^{n-1}(6)+2)/(2n+1+(-1)^{n-1})\]

OpenStudy (inkyvoyd):

THat's undefined

OpenStudy (inkyvoyd):

Infinity isn't necessarily even or odd

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac{(-1)^n6+2}{2n+1+(-1)^n}\] like that ?

OpenStudy (anonymous):

no, satellite to the power of n-1

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac{(-1)^n6+2}{2n+1+(-1)^{n-1}}\] like this?

OpenStudy (anonymous):

there is no "n" in the numerator, so as n goes to infinity so does the denominator. the numerator is either -4 or 8 but in any case the whole thing goes to zero

OpenStudy (anonymous):

can I assume the limit to infinity of the absolute value goes to zero? or does taking the absolute value break the alternating?